# Non-Hermitian quantum geometric tensor and nonlinear electrical response

### v1

@@ Abstract
Effect of Wavepacket Width on SODCc. Assuming the imaginary component of the band dispersion is perturbatively small compared to all other energy scales (e.g., real band energies and scattering rates), we derive the SODCc as (see Appendix A for details):
$\sigma_{\mu\nu\theta}^{W}=\sigma_{\mu\nu\theta}^{W=0}+e^{3}\int_{\mathbf{k}}W^{2}\tau\Gamma_{2,1}+W^{2}\tau^{2}\Gamma_{2,2}+W^{4}\tau^{2}\Gamma_{4,2},$ (6)
@@
The term $\sigma_{\mu\nu\theta}^{W=0}$ represents the DC conductivity in the narrow wavepacket width limit, i.e., Eq. (1). The coefficients $\Gamma_{2,1}$, $\Gamma_{2,2}$, and $\Gamma_{4,2}$ encode the interplay between finite wavepacket width and the complex QGT. These terms are defined as
@@
$\Gamma_{4,2}$ $=\frac{\epsilon_{i\nu j}\epsilon_{n\theta\gamma}\,\mathrm{Im}\Omega_{j}\,\partial_{i}\left(\mathrm{Im}\Omega_{\gamma}\partial_{n}f_{0}\right)+(\theta\leftrightarrow\nu)}{2}\,\partial_{\mu}\mathrm{Re}\xi_{\mathbf{k}}^{(0)}.$
The coefficients $\Gamma_{2,2}$ and $\Gamma_{4,2}$ depend solely on the imaginary part of the non-Hermitian Berry curvature. In contrast, the coefficient $\Gamma_{2,1}$ involves contributions from both the full complex non-Hermitian Berry curvature and the imaginary part of the non-Hermitian band-renormalized QM.
To elucidate the role of the wavepacket width in the SODCc, we introduce a small imaginary perturbation to the parameter $m$ in the model defined by the Hamiltonian (5). This modification results in a Hamiltonian with complex eigenvalues and a line gap [Fig. 4(a)], while the QM and Berry curvature acquire complex character, as illustrated in Figs. 4(c–f). Consequently, all coefficients $\Gamma_{i,j}$ become non-zero, leading to a wavepacket-width-dependent enhancement of $\sigma_{xxx}^{W}$, as demonstrated in Fig. 4(b).
@@
Conclusions and Outlook. In this work, we reveal the role of the QGT in governing nonlinear electrical responses in non-Hermitian systems with a spectral line gap. In the narrow-wavepacket-width limit, the SODCc depends solely on the real part of the band dispersion, the Berry curvature, and the band-renormalized QM. Beyond this limit, the wavepacket width bridges the complex QM and Berry curvature—unique features of the projected Hilbert space in non-Hermitian systems—to the nonlinear response. Using 1D and 2D non-Hermitian models, we demonstrate the distinct role of the QM in shaping the SODCc. Notably, in the narrow-wavepacket regime, our results suggest that the band-renormalized QM governs the SODCc independently of the scattering rate. Furthermore, the wavepacket width amplifies the SODCc, a phenomenon unique to non-Hermitian systems arising from the imaginary components of the band dispersion and Berry curvature. These contributions introduce terms proportional to the wavepacket width that modify the semiclassical equations of motion.
@@
Our results extend to surfaces and edges of topological materials, such as Chern insulators, which are effectively described by non-Hermitian Hamiltonians. This connection enables predictions of surface nonlinear conductivity in Hermitian systems. By bridging quantum geometry with nonlinear transport, this work establishes a framework to harness geometric effects in topological devices and engineered materials.
@@
Acknowledgments K.C. acknowledges support from the Fundamental Research Funds for the Central Universities. J.Z. acknowledges support from the National Natural Science Foundation of China (Grants No. 92263208), the National Key R&D Program of China (Grants No. 2022YFA1404400), and the Research Grants Council of Hong Kong SAR (Grant No. AoE/P-502/20).
χ: KaiChenPhys@tongji.edu.cn +: jiezhu@tongji.edu.cn
@@ Appendix A Nonlinear electric response and the quantum geometric tensor
$f_{0}=\frac{1}{1+\exp\left(\mathrm{Re}\,\xi_{\mathbf{k}}/k_{B}T\right)},$
with the Boltzmann constant $k_{B}$ set to 1 throughout this work.
@@
where $\mathbf{W}^{\left(0\right)}_{\mathbf{k}}=W^{2}\nabla_{\mathbf{k}}\text{Im}\xi^{(0)}_{\mathbf{k}}$, $\mathbf{W}^{\left(1\right)}_{\mathbf{k}}=-e\mathbf{E}+W^{2}\text{Im}\left[\nabla_{\mathbf{k}}\xi^{(1)}_{\mathbf{k}}+e\mathbf{E}\times\bm{\Omega}^{(0)}_{\mathbf{k}}\right$] and $\mathbf{W}^{\left(2\right)}_{\mathbf{k}}=W^{2}\text{Im}\left[\nabla_{\mathbf{k}}\xi^{(2)}_{\mathbf{k}}+e\mathbf{E}\times\bm{\Omega}^{(1)}_{\mathbf{k}}\right$]. In the direct current limit, where $\omega\ll 1/\tau$, the above equations simplify to:
@@
A.1 narrow wave packet width
In this subsection, we consider the limit of a narrow wave packet width, such that the term proportional to $W^{2}$ in $\mathbf{W}_{\mathbf{k}}$ can be safely neglected. We then obtain:
@@
where, in the first equation, we choose a gauge for the Berry curvature to make it perpendicular to the electric field. Here, $G_{n,\mu\nu}^{LR}=\sum_{i\neq n}\frac{A^{LR}_{ni,\mu}A^{LR}_{in,\nu}+A^{LR}_{ni,\nu}A^{LR}_{in,\mu}}{2\left(\xi^{(0)}_{n,\mathbf{k}}-\xi^{(0)}_{i,\mathbf{k}}\right)}$ denotes the band-renormalized non-Hermitian quantum metric.
@@ A.2 Effects of Wave Packet Width
In the previous subsection, we considered the limit of a narrow wave packet, where the time evolution of momentum is solely dictated by the external electric field. However, in this limit, certain non-Hermitian effects are suppressed. Here, we extend our analysis to finite wave packet widths and examine their impact on the DC conductivity in non-Hermitian systems. To facilitate analytical progress, we focus on a non-Hermitian model in which the imaginary part of the spectrum varies slightly in $\mathbf{k}$-space within certain parameter regimes. In this case, the semiclassical equations of motion are modified as follows:
@@
$+\tau^{2}e^{3}W^{4}\int_{\mathbf{k}}\frac{\epsilon_{i\nu j}\epsilon_{n\theta\gamma}\text{Im}\Omega_{j}\partial_{i}\left(\text{Im}\Omega_{\gamma}\partial_{n}f_{0}\right)+\theta\leftrightarrow\nu}{2}\partial_{\mu}\text{Re}\xi_{\mathbf{k}}^{(0)}$
@@
$+\tau^{2}e^{3}W^{4}\int_{\mathbf{k}}\frac{\epsilon_{i\nu z}\epsilon_{n\theta z}\text{Im}\Omega_{z}\partial_{i}\left(\text{Im}\Omega_{z}\partial_{n}f_{0}\right)+\theta\leftrightarrow\nu}{2}\partial_{\mu}\text{Re}\xi_{\mathbf{k}}^{(0)}$
@@
$+\tau^{2}e^{3}W^{4}\int_{\mathbf{k}}\epsilon_{i\mu z}\epsilon_{n\mu z}\text{Im}\Omega_{z}\partial_{i}\left(\text{Im}\Omega_{z}\partial_{n}f_{0}\right)\partial_{\mu}\text{Re}\xi_{\mathbf{k}}^{(0)}$
@@
$+\tau^{2}e^{3}W^{4}\int_{\mathbf{k}}\text{Im}\Omega_{z}\partial_{y}\left(\text{Im}\Omega_{z}\partial_{y}f_{0}\right)\partial_{x}\text{Re}\xi_{\mathbf{k}}^{(0)}$
@@
$+\tau^{2}e^{3}W^{4}\int_{\mathbf{k}}\text{Im}\Omega_{z}\partial_{x}\left(\text{Im}\Omega_{z}\partial_{x}f_{0}\right)\partial_{y}\text{Re}\xi_{\mathbf{k}}^{(0)}$

### v2

@@ II Nonlinear DC conductivity and QGT in non-Hermitian systems
In the previous section, we derived the SODCc in the limit of zero wave-packet width. We now extend the analysis to incorporate a finite wave-packet width, which introduces novel contributions governed by the complex structure of the QGT. Assuming the imaginary component of the band dispersion is perturbatively small compared to all other energy scales (e.g., real band energies and scattering rates), we derive the SODCc as (see Appendix A for details):
$\sigma_{\mu\nu\theta}^{W}=\sigma_{\mu\nu\theta}^{W=0}+\int_{\mathbf{k}}W^{2}\tau\Gamma_{2,1}+W^{2}\tau^{2}\Gamma_{2,2}+W^{4}\tau^{2}\Gamma_{4,2},$ (9)
@@
The term $\sigma_{\mu\nu\theta}^{W=0}$ represents the DC conductivity in the zero wavepacket width limit, i.e., Eq. (7). The coefficients $\Gamma_{2,1}$, $\Gamma_{2,2}$, and $\Gamma_{4,2}$ encode the interplay between finite wavepacket width and the complex QGT. These terms are defined as
@@
$\Gamma_{4,2}$ $=\frac{\epsilon_{i\nu j}\epsilon_{n\theta\gamma}\,\mathrm{Im}\Omega^{(0)}_{j}\,\partial_{i}\left(\mathrm{Im}\Omega^{(0)}_{\gamma}\partial_{n}f_{0}\right)+(\theta\leftrightarrow\nu)}{2}\,\partial_{\mu}\mathrm{Re}\xi_{\mathbf{k}}^{(0)}.$
Having presented the general result, we can now analyze its geometric content. The coefficients $\Gamma_{2,2}$ and $\Gamma_{4,2}$ depend solely on the imaginary part of the non-Hermitian Berry curvature. In contrast, the coefficient $\Gamma_{2,1}$ involves contributions from both the full complex non-Hermitian Berry curvature and the imaginary part of the non-Hermitian band-renormalized QM. This structure reveals a key physical insight: the finite wavepacket width couples directly to the geometric quantities. This coupling generates distinct conductivity contributions scaling with $W^{2}$ and $W^{4}$, which are unique signatures of non-Hermitian quantum geometry and have no counterpart in Hermitian systems.
Having established the theoretical result, we now address its experimental implications. A natural question concerns the experimental relevance of the wavepacket width $W$, which governs the magnitude of the geometric nonlinear response (scaling as $W^{2}$). In solid-state quantum materials, effective non-Hermiticity naturally arises from finite quasiparticle lifetimes or coupling to an environment. In this regime, the effective spatial extent of the wavepacket $W$ is intrinsic and is typically determined by the carrier coherence length. Semiclassically, this width is physically associated with the thermal de Broglie wavelength, $\lambda_{th}\propto T^{-1/2}$ [17, 64].
Consequently, the parameter $W$ is naturally tunable via temperature. This suggests that the geometric contributions to the nonlinear conductivity (proportional to $W^{2}$) will exhibit a distinct $T^{-1}$ scaling. This specific temperature dependence allows the geometric non-Hermitian response to be experimentally disentangled from extrinsic scattering-induced mechanisms. Conversely, this relationship implies that nonlinear transport measurements can serve as a diagnostic tool: by isolating the $T^{-1}$ component of the nonlinear conductivity, one can experimentally extract the effective coherence length of the non-Hermitian system.
IV SODCc in one-dimensional non-Hermitian models.
To illustrate how the QGT manifests in a non-Hermitian system’s SODCc, in this section we consider two concrete models within the zero-width wavepacket approximation: a paradigmatic one-dimensional lattice model and a more complex boundary system. The effect of a finite wavepacket width on the SODCc will be examined for a two-dimensional model in the next section.
IV.1 Non-Hermitian Su-Schrieffer-Heeger model.
Fig. 1: (Color online) Schematic of the non-Hermitian SSH lattice model.
We begin with the non-Hermitian Su-Schrieffer-Heeger (SSH) model, a one-dimensional topological insulator featuring nonreciprocal intracell hopping [65]. The lattice structure is schematically depicted in Fig. 1. The Bloch Hamiltonian is given by
$H_{\text{nHSSH}}(k)$ $=\left(\frac{t_{1}+t_{2}}{2}+t\cos k\right)\sigma_{x}$ (11)
$\quad+\left(i\frac{t_{1}-t_{2}}{2}+t\sin k\right)\sigma_{y},$
where $\sigma_{i}$ ($i=x,y$) are the Pauli matrices, $t\in\mathbb{R}$ is the intercell hopping, and $t_{1}\in\mathbb{C}$ and $t_{2}\in\mathbb{R}$ are the intracell directional hoppings. The Hamiltonian becomes Hermitian when $t_{1}=t_{2}\in\mathbb{R}$.
For this one-dimensional model, the Berry curvature vanishes. Consequently, the general expression for the SODCc in Eq. (7) simplifies significantly to:
$\sigma_{xxx}=-\text{Re}\int_{\mathbf{k}}f_{0}\left\{\tau^{2}\partial^{3}_{x}\xi^{(0)}_{\mathbf{k}}+\partial_{x}G^{LR}_{xx}\right\}.$ (12)
This simplified form isolates two contributions: a conventional Drude-like term ($\propto\tau^{2}$) and an intrinsic geometric term involving the gradient of the QM. The non-Hermitian SSH model preserves time-reversal symmetry, given by $H_{\text{nHSSH}}(k)^{*}=H_{\text{nHSSH}}(-k)$ when $t_{1}\in\mathbb{R}$. Under this symmetry, the band dispersion satisfies $\xi^{(0)*}_{\mathbf{k}}=\xi^{(0)}_{-\mathbf{k}}$, and the Berry connection obeys $A^{LR(0)}_{mn,\nu}(k)^{*}=A^{LR(0)}_{mn,\nu}(-k)$, which forces $\sigma_{xxx}=0$. Therefore, we introduce a complex $t_{1}$ to break time-reversal symmetry, while carefully choosing its value to maintain a finite line gap [Fig. 2(a)].
We now analyze the resulting conductivity. As shown in Eq. (12), the real part of the band-renormalized QM provides an intrinsic contribution to the SODCc, which is independent of $\tau$. Its temperature dependence is shown in Fig. 2(b). The ratio of the intrinsic term $\sigma_{xxx}^{0}$ to the total SODCc $\sigma_{xxx}$ increases with $T$. In the strong scattering regime ($\tau=1$), the intrinsic term dominates the SODCc [Fig. 2(c)]. In contrast, in the weak scattering regime ($\tau=10$), the intrinsic contribution is small at low $T$ but increases with temperature.
The behavior near the topological phase transition is particularly revealing. The energy gap of the non-Hermitian SSH model closes at the critical point $t_{2}/t=1$. Near this critical value, the intrinsic contribution dominates the SODCc in both the strong ($\tau=1$) and weak ($\tau=10$) scattering regimes [Fig. 2(d)]. This demonstrates that quantum geometric effects become particularly pronounced near topological phase transitions, where the band-renormalized quantum metric becomes singular.
IV.2 Boundary of a Chern insulator.
Having examined a minimal lattice model, we now consider a system with a more intricate physical origin: the effective boundary of a Chern insulator. This example illustrates how non-Hermitian Hamiltonians emerge naturally in subsystems coupled to an environment. Previous studies [66] have shown that the low-energy physics of this boundary can be effectively captured by a non-Hermitian Hamiltonian, where non-Hermiticity arises from the coupling between the boundary and the bulk, thereby inducing a complex self-energy. The effective non-Hermitian Hamiltonian takes the form
This model also exhibits a phase with a line gap [see Fig. 3(a)]. We compute its SODCc to test the generality of the geometric effects found in the SSH model. As illustrated in Fig. 3(b), the intrinsic contribution to the total SODCc increases with temperature in both the strong and weak scattering regimes. At low temperatures, similar to the non-Hermitian SSH model, the intrinsic contribution is more pronounced in the strong scattering regime than in the weak scattering regime.
Together, these two examples illustrate the applicability of our framework to distinct physical settings. The non-Hermitian SSH model shows how intrinsic geometric conductivity arises in a simple, tunable lattice. The Chern insulator boundary model confirms that these effects persist in more complex, physically motivated systems where non-Hermiticity emerges from environmental coupling. In both cases, the QGT provides the fundamental link between the complex band structure and the measurable nonlinear response.
V SODCc in a two-dimensional non-Hermitian model and the effect of wavepacket width.
We now extend our analysis to a two-dimensional system to investigate the effect of a finite wavepacket width on the SODCc. To isolate this effect, we first examine a model with a real energy spectrum, where the QGT components are purely real. We then introduce a controlled perturbation to generate complex eigenvalues and complex QGT components, enabling us to study the direct coupling between wavepacket width and the imaginary parts of the geometry.
We consider a two-dimensional non-Hermitian model described by the Hamiltonian [67]:
where the function $h(k_{x},k_{y})=e^{-ik_{y}}+2e^{ik_{y}/2}\cos\left(\frac{\sqrt{3}}{2}k_{x}\right)$ defines the system describing a non-reciprocal lattice such that for $\gamma\neq 1$ the Hamiltonian is non-Hermitian yet can exhibit real eigenvalues [Fig. 4(a)]. Note that the reality of the spectrum in Eq. 14 arises from its pseudo-Hermitian construction [67], which guarantees real eigenvalues for all $\gamma>0$, distinct from $\mathcal{PT}$-symmetric systems that generally undergo symmetry breaking.
The real eigenvalues of the Hamiltonian allow us to disentangle the effects of the non-Hermitian QGT from the dissipative or amplifying effects typically associated with complex eigenvalues. Furthermore, both the Berry curvature and the QM are real-valued functions of the Bloch momentum $\mathbf{k}$; accordingly, Figs. 4(b) and 4(d) display the distributions of the Berry curvature $\text{Re}\Omega_{z}(k_{x},k_{y})$ and the QM component $\text{Re}G_{xy}$ in $\mathbf{k}$-space. Notably, the Berry curvature satisfies $\Omega_{z}(k_{x},k_{y})=-\Omega_{z}(-k_{x},k_{y})$ and $\Omega_{z}(k_{x},k_{y})=\Omega_{z}(k_{x},-k_{y})$, while the QM component obeys $G_{xy}(k_{x},k_{y})=-G_{xy}(-k_{x},k_{y})$ and $G_{xy}(k_{x},k_{y})=-G_{xy}(k_{x},-k_{y})$. Taking the xyyxyy-component of the SODCc, $\sigma_{xyy}$, as a concrete example and noting that $E_{\pm}(\mathbf{k})=E_{\pm}(-\mathbf{k})$, both the first- and second-order $\tau$ terms in Eq. (12) vanish, so that only the band-renormalized QM, which is independent of $\tau$, contributes to $\sigma_{xyy}$ [see Fig. 4(c)].
Having established the behavior for the purely real-spectrum case, we now examine the effect of finite wavepacket width on the SODCc. In non-Hermitian systems, the QGT possesses a complex structure. Whereas the imaginary part of the QGT in Hermitian systems corresponds to the purely real Berry curvature, the Berry curvature in non-Hermitian systems becomes intrinsically complex. This complexity introduces an imaginary component that couples dynamically to the spatial width of quantum wavepackets via the semiclassical equation $\dot{\mathbf{k}}=-e\mathbf{E}+W^{2}\,\text{Im}\!\left[\nabla_{\mathbf{k}}\xi_{n,\mathbf{k}}+e\mathbf{E}\times\bm{\Omega}_{n,\mathbf{k}}\right]$. The conventional zero-width approximation suppresses this coupling by neglecting finite-$W$ contributions. To reveal the role of this coupling in the nonlinear response, we introduce a small imaginary perturbation to the parameter $m$ in the Hamiltonian (14). This yields a system with complex eigenvalues and a line gap [Fig. 5(a)], while the QM and Berry curvature acquire a complex character [Figs. 5(c–f)]. As a result, the coefficients $\Gamma_{i,j}$ become nonzero, and a finite wavepacket width produces a pronounced enhancement of $\sigma_{xxx}^{W}$, as shown in Fig. 5(b).
VI Conclusions and Outlook.
@@
In this work, we have elucidated the role of the QGT in governing nonlinear electrical responses in non-Hermitian systems with a spectral line gap. In the zero-wavepacket-width limit, the SODCc depends on the real part of the band dispersion, the Berry curvature, and the band-renormalized QM. Beyond this limit, the finite wavepacket width bridges the complex-valued band-renormalized QM and Berry curvature—unique features of the projected Hilbert space in non-Hermitian systems—to the nonlinear response. Using one- and two-dimensional non-Hermitian models, we have demonstrated the distinct role of the QM in shaping the SODCc. Notably, in the zero-width regime, our results indicate that the band-renormalized QM dominates the SODCc at higher temperatures and under conditions of strong scattering (short $\tau$). Furthermore, a finite wavepacket width amplifies the SODCc, a phenomenon exclusive to non-Hermitian systems that arises from the imaginary components of the band dispersion and Berry curvature.
@@
Our findings extend naturally to surfaces and edges of topological materials, such as Chern insulators, which are often effectively described by non-Hermitian Hamiltonians. This connection enables the prediction of surface nonlinear conductivity in otherwise Hermitian systems. By bridging quantum geometry with nonlinear transport, this work establishes a versatile framework for harnessing geometric effects in topological devices and engineered materials.
Acknowledgements.
@@
Acknowledgments—K.C. acknowledges support from the Shanghai Magnolia (Yulan) Pu River Project (Grant No. 24PJD121) and from the Fundamental Research Funds for the Central Universities. J.Z. acknowledges support from the National Natural Science Foundation of China (Grants No. 92263208), the National Key R&D Program of China (Grants No. 2022YFA1404400), and the Research Grants Council of Hong Kong SAR (Grant No. AoE/P-502/20). χ: KaiChenPhys@tongji.edu.cn +: jiezhu@tongji.edu.cn
@@ Appendix A Nonlinear electric response and the quantum geometric tensor
$f_{0}=\frac{1}{1+\exp\left(\mathrm{Re}\,\xi_{\mathbf{k}}/k_{B}T\right)}.$
@@
where $\mathbf{W}^{\left(0\right)}_{\mathbf{k}}=W^{2}\nabla_{\mathbf{k}}\text{Im}\xi^{(0)}_{\mathbf{k}}$, $\mathbf{W}^{\left(1\right)}_{\mathbf{k}}=-e\mathbf{E}+W^{2}\text{Im}\left[\nabla_{\mathbf{k}}\xi^{(1)}_{\mathbf{k}}+e\mathbf{E}\times\bm{\Omega}^{(0)}_{\mathbf{k}}\right$] and $\mathbf{W}^{\left(m\geq 2\right)}_{\mathbf{k}}=W^{2}\text{Im}\left[\nabla_{\mathbf{k}}\xi^{(m)}_{\mathbf{k}}+e\mathbf{E}\times\bm{\Omega}^{(m-1)}_{\mathbf{k}}\right$]. In the direct current limit, where $\omega\ll 1/\tau$, the above equations simplify to:
@@
A.1 zero wavepacket width
In this subsection, we consider the limit of a zero wave packet width, such that the term proportional to $W^{2}$ in $\mathbf{W}_{\mathbf{k}}$ can be safely neglected. We then obtain:
@@
where, in the first equation, we choose a gauge for the Berry curvature to make it perpendicular to the electric field. Here, $G_{n,\mu\nu}^{LR}=\sum_{i\neq n}\frac{A^{LR}_{ni,\mu}A^{LR}_{in,\nu}+A^{LR}_{ni,\nu}A^{LR}_{in,\mu}}{2\left(\xi^{(0)}_{n,\mathbf{k}}-\xi^{(0)}_{i,\mathbf{k}}\right)}$ denotes the band-renormalized non-Hermitian QM.
@@ A.2 Effects of Wave Packet Width
In the previous subsection, we considered the limit of a zero wave packet, where the time evolution of momentum is solely dictated by the external electric field. However, in this limit, certain non-Hermitian effects are suppressed. Here, we extend our analysis to finite wave packet widths and examine their impact on the DC conductivity in non-Hermitian systems. To facilitate analytical progress, we focus on a non-Hermitian model in which the imaginary part of the spectrum varies slightly in $\mathbf{k}$-space within certain parameter regimes. In this case, the semiclassical equations of motion are modified as follows:
@@
$+\tau^{2}e^{3}W^{4}\int_{\mathbf{k}}\frac{\epsilon_{i\nu j}\epsilon_{n\theta\gamma}\text{Im}\Omega^{(0)}_{j}\partial_{i}\left(\text{Im}\Omega^{(0)}_{\gamma}\partial_{n}f_{0}\right)+\theta\leftrightarrow\nu}{2}\partial_{\mu}\text{Re}\xi_{\mathbf{k}}^{(0)}$
@@
$+\tau^{2}e^{3}W^{4}\int_{\mathbf{k}}\frac{\epsilon_{i\nu z}\epsilon_{n\theta z}\text{Im}\Omega^{(0)}_{z}\partial_{i}\left(\text{Im}\Omega^{(0)}_{z}\partial_{n}f_{0}\right)+\theta\leftrightarrow\nu}{2}\partial_{\mu}\text{Re}\xi_{\mathbf{k}}^{(0)}$
@@
$+\tau^{2}e^{3}W^{4}\int_{\mathbf{k}}\epsilon_{i\mu z}\epsilon_{n\mu z}\text{Im}\Omega^{(0)}_{z}\partial_{i}\left(\text{Im}\Omega^{(0)}_{z}\partial_{n}f_{0}\right)\partial_{\mu}\text{Re}\xi_{\mathbf{k}}^{(0)}$
@@
$+\tau^{2}e^{3}W^{4}\int_{\mathbf{k}}\text{Im}\Omega^{(0)}_{z}\partial_{y}\left(\text{Im}\Omega^{(0)}_{z}\partial_{y}f_{0}\right)\partial_{x}\text{Re}\xi_{\mathbf{k}}^{(0)}$
@@
$+\tau^{2}e^{3}W^{4}\int_{\mathbf{k}}\text{Im}\Omega^{(0)}_{z}\partial_{x}\left(\text{Im}\Omega^{(0)}_{z}\partial_{x}f_{0}\right)\partial_{y}\text{Re}\xi_{\mathbf{k}}^{(0)}$